. (a) Calculate the self-inductance of a toroidal coil consisting of 1000 turns of wire, each of radius 1 cm wound uniformly on a non-ferromagnetic ring of mean radius 10 cm. (10)


(b) A circuit containing an air-cored solenoid of self-inductance 0.1 H carries a steady current of 1 A. A switch in the circuit is opened, reducing the current to zero in a time of 1 ms. Determine the magnitude of the induced emf across the solenoid, assuming the current fallsto zero at a constant rate.


1
Expert's answer
2021-06-30T10:19:12-0400

(a)

"L = \frac{\mu_0 N^2A}{2γ} \\

\mu_0 = 4 \pi \times 10^{-7}\;H/m \\

Area=A=\pi r^2 \\

r=radius = 10 \;cm = 0.1 \;m \\

N= number \;of \;turns = 1000 \\

L = \frac{4 \pi \times 10^{-7} \times (1000)^2 \times (0.1)^2}{2 \times 0.1} \\

L=0.062\;H"

(b)

Initial current

I1 = 1 A

Final current

I2=0

Time interval

t = 1ms "= 10^{-3} \;s"

Self-inductance of a solenoid

"L= \frac{\mu_0N^2A}{l}"

Magnetic flux

"Ø = BA \\

= \frac{\mu_0NI}{l}A \\

= \frac{\mu_0N^2A}{l} \times \frac{I}{N} \\

Ø = L \times \frac{I}{N} \\

E_{ind}=-N\frac{ΔØ}{Δt} \\

= -N \frac{L}{N} \times \frac{ΔI}{Δt} \\

= - \frac{L(I_2-I_1)}{Δt} \\

E_{ind} = - \frac{0.1(0-2)}{10^{-3}} = 100 \;V"